Project

General

Profile

2019-03-12-MEMOMIPO

Joint MICE Experiment Management Office and MICE International Project Office

12th March 2019; 15:30 GMT

Phone details

Agenda

  1. Introduction: KL
  2. Notes of previous meeting and status of actions: All
    • KL: migrate CB mailing lists to jiscmail
      • Done: CB list now done, contents being checked by PK.
    • CR: migrate analysis mailing list to jiscmail
    • KL: Doodle for next amplitude evolution paper referees meeting.
      • Done.
    • PK: Work with CB to complete Grey Book entry for MICE.
  3. Software and computing status: PF
  4. Paper status and plans: CR
  5. AOB

Notes

Present: KL, CR, PK, SB, PF, CW
Apologies:

  1. Introduction: KL
  2. Notes of previous meeting and status of actions: All
    • KL: migrate CB mailing lists to jiscmail
      • Done: CB list now done, contents being checked by PK.
    • CR: migrate analysis mailing list to jiscmail.
      • Stands.
    • KL: Doodle for next amplitude evolution paper referees meeting.
      • Done.
    • PK: Work with CB to complete Grey Book entry for MICE.
      • In progress: PK plans to implement Brunel so that he knows the system. Then he will ask CB reps to send him the relevant form so that he can liaise with CERN.
  3. Software and computing status: PF
    • LFC workaround still being implemented. J.Martyniak will migrate the workaround he is coding for SHiP for MICE.
    • MAUS 3.3 is ready on the server. Will begin to use it soon.
    • PF is checking details of the geometry being used by CR that has been tuned to reproduce the track-chi2 distribution etc.
    • PK has downloaded and compiled MAUS as part of the handover from PF.
  4. Paper status and plans: CR
    • LiH scattering: work continues, no update at this meeting.
    • LH2 scattering: GI has observed 'alignment' mismatches between tracks u/s and d/s of the absorber when they are migrated back to the absorber. Now investigating. Could it be that the alignment prepared by F.Drielsma does not work for some subset of runs. Under investigation.
    • Field on scattering:
    • Magnetic field allignment: J.Langlands will leave Sheffield in mid April. Therefore, we carry a risk that the magnetic field issues may persist once he has left.
    • Emittance measurement: V.Blackmore has sent corrections to the proofs to EJPC. We await publication.
    • Amplitude evolution: work progresses following feedback at CM53 and most recent referees' meeting.
    • CR had prepared a note describing the b/l for distribution to possible future users. Comments were solicited.
    • Discussion:
      • CW asked what the status of IPAC material is for his posters. CR reminded CW of the rules; indicated new material would need to be presented at April 18 VC and papers/posters need to be distributed to referees two weeks before the conference.
      • PF commented on the beamline note that CR uploaded; that the target would need to be checked. We noted that other aspects of the hardware may also need verification before taking on a new project using the MICE beam.
  5. AOB

Summary of actions:

Back to MEMO:

Back

Back to MIPO:

Back

Documents

MICE_Beamline_Description.pdf (525 KB) Rogers, Chris, 12 March 2019 15:34